### v2

@@ Abstract
With the global population growing and arable land resources becoming increasingly scarce, smart agriculture and precision agriculture have emerged as key directions for the future of agricultural development. Artificial intelligence (AI) technologies, particularly deep learning models, have found widespread applications in areas such as crop monitoring and pest detection. As an emerging generative model, diffusion models have shown significant promise in tasks like agricultural image processing, data augmentation, and remote sensing. Compared to traditional generative adversarial networks (GANs), diffusion models offer superior training stability and generation quality, effectively addressing challenges such as limited agricultural data and imbalanced image samples. This paper reviews the latest advancements in the application of diffusion models in agriculture, focusing on their potential in crop pest and disease detection, remote sensing image enhancement, crop growth prediction, and agricultural resource management. Experimental results demonstrate that diffusion models significantly improve model accuracy and robustness in data augmentation, image generation, and denoising, especially in complex environments. Despite challenges related to computational efficiency and generalization capabilities, diffusion models are expected to play an increasingly important role in smart and precision agriculture as technology advances, providing substantial support for the sustainable development of global agriculture.
@@
diffusion model, smart agriculture, precision agriculture, pest and disease detection, data augmentation, image generation, remote sensing images, crop monitoring, deep learning.
@@ 1 Introduction
With the continuous growth of the global population and the increasing scarcity of arable land resources, traditional agriculture is facing unprecedented challenges. To improve agricultural production efficiency, ensure food security, and achieve sustainable development, countries worldwide are actively advancing the modernization of agriculture. In this context, smart agriculture and precision agriculture, as key components of agricultural modernization, are gradually becoming the mainstream directions for future agricultural development.
@@
Smart agriculture leverages modern information technologies such as artificial intelligence (AI), the Internet of Things (IoT), big data, and remote sensing to enable dynamic sensing and data-driven decision-making regarding key factors such as crop growth environments, agricultural machinery operational status, soil moisture, and climate change. By enabling precise fertilization, targeted irrigation, and intelligent pest and disease identification, smart agriculture not only significantly improves resource utilization but also enhances the intelligence, automation, and sustainability of agricultural production.As artificial intelligence technology continues to evolve, deep learning models, especially image generation models, have become a focal point in agricultural intelligence research. Among these, diffusion models, a new and advanced type of generative model in recent years, have demonstrated capabilities far exceeding traditional methods in tasks such as image generation [1, 2, 3, 4, 5, 6, 7], image restoration [1, 3, 4, 8, 9, 10], image segmentation [11, 12, 13, 14], and image enhancement [15]. They have achieved significant results in fields such as medicine [11, 13], industrial inspection [16], and artistic creation. Introducing these models into agricultural scenarios provides new solutions to real-world challenges, such as limited agricultural data, imbalanced samples, and insufficient data diversity. Therefore, systematically reviewing the current state of research and typical applications of diffusion models in agriculture can help promote the deep integration of AI technology with modern agriculture and accelerate the practical application of agricultural intelligence.In the process of promoting agricultural intelligence, deep learning, with its powerful feature extraction and pattern recognition capabilities, has become an essential technological support for agricultural informatization and automation. Compared to traditional machine learning methods, deep learning can automatically learn complex nonlinear relationships through an end-to-end approach, significantly enhancing the analysis of images, time series, and multi-source data. It is widely applied in various agricultural subfields, such as crop identification, pest and disease detection [17, 18], yield prediction [19], and environmental monitoring [20, 21, 22].
@@
In image processing, convolutional neural networks (CNNs) [23] are widely used for tasks such as crop classification [24], weed identification, and leaf disease diagnosis [17, 18], leveraging high-resolution images to identify minute differences in agricultural fields. In remote sensing agriculture, deep neural networks can effectively process remote sensing data such as radar [25, 26], and multispectral images [27, 28, 29], enabling the automatic identification and analysis of crop types, planting areas, growth conditions, and moisture status. In time series prediction, recurrent neural networks (RNNs) and their variants, such as long short-term memory (LSTM) networks, are employed to model agricultural time series data like soil moisture, meteorological parameters [22], and crop growth indicators, providing valuable data for precise fertilization and irrigation decisions.
@@
Additionally, generative models such as generative adversarial networks (GANs) [15, 30] are being increasingly used for agricultural data synthesis and augmentation to address sample scarcity and data imbalance in real-world agricultural scenarios. However, GANs suffer from issues like training instability and model collapse, which limit their further application in agriculture. Against this backdrop, the stability and high-quality generation capabilities of diffusion models are increasingly demonstrating their potential value in agricultural applications.
@@
Diffusion models are a class of deep generative models that have made breakthrough progress in generative modeling in recent years. Their core concept originates from the mathematical modeling of diffusion processes in non-equilibrium thermodynamics. Unlike generative adversarial networks (GANs) [15, 30] and variational autoencoders (VAEs) [31, 32], diffusion models generate high-quality samples by gradually adding noise to the data during the training phase (the forward process) and then gradually denoising it during the generation phase using the learned reverse process. Since their introduction, diffusion models have undergone significant development and achieved breakthroughs across various fields. The concept of diffusion models originally emerged from diffusion processes in physics, aiming to simulate the gradual evolution of data from a true distribution to a noisy distribution and then recover the data from noise through the inverse process. In 2020, Ho et al. proposed Denoising Diffusion Probabilistic Models (DDPM), marking the first large-scale application of diffusion models in deep learning and addressing the issue of training instability in generative models, particularly in the field of image generation. The introduction of DDIM further optimized generation efficiency by reducing the number of generation steps through an implicit diffusion process, thus improving generation speed. Meanwhile, Score Matching Generative Models (SGM) introduced a new generation approach for diffusion models by learning the score function of data for generation, thus avoiding the instability issues present in traditional generative models. As the technology evolved, conditional diffusion models were proposed to handle more flexible conditional generation tasks, such as text-to-image generation, expanding the application scenarios of diffusion models. In 2022, latent diffusion models (LDMs) significantly improved generation efficiency and reduced computational requirements by performing diffusion in the latent space, further promoting the widespread application of diffusion models, especially in high-resolution image generation and artistic creation. The introduction of Stable Diffusion marked a new era for diffusion models, becoming an important tool in the field of text-to-image generation due to its efficient generation capabilities and open-source nature. Although diffusion models excel in terms of generation quality and training stability, they still face challenges such as computational resource demands, dataset quality, and generalization capabilities. In the future, with improvements in computational efficiency, the development of multi-modal data fusion technologies, and model optimization in low-resource environments, diffusion models will play a significant role in a broader range of application scenarios, driving the further development of generative artificial intelligence.
The overall structure of this paper is as follows: Section 1 introduces the challenges and technological needs facing agriculture, as well as the potential, advantages, and challenges of diffusion models in agriculture. Section 2 covers the theory, evolution, and variants of diffusion models, as well as their relationship with other generative models. Section 3 discusses the specific applications of diffusion models in image processing in the agricultural field, including agricultural image classification, data augmentation, image generation, and pest and disease diagnosis. Section 4 compares diffusion models with other methods and conducts experiments and analyses. Section 5 explores the future prospects of diffusion models in the field of agricultural images. Section 6 summarizes the content of the paper.
The contributions of this paper are as follows: (1) We systematically reviewed seven typical application areas of diffusion models in smart agriculture (as shown in Figure 10), marking the first comprehensive review in this field; (2) We compared and analyzed the performance differences between diffusion models and GANs in agricultural imaging tasks, conducting the first small-scale experimental validation of the advantages of diffusion models in data augmentation for long-tail distributions; (3) We summarized the key challenges currently limiting the large-scale application of diffusion models (such as computational efficiency, data bottlenecks, and model generalization) and outlined future research directions.
@@ 2.1 Major Challenges and Technological Needs Facing Agriculture
The primary challenges facing agriculture include crop health and pest management, precision agriculture and resource management, as well as remote sensing image processing and data analysis. Pests and diseases are key factors affecting agricultural production. Traditional manual inspection methods are inefficient and often lead to misdiagnosis, particularly in the early stages of disease. As agricultural scales expand, intelligent technologies have become essential for improving the efficiency and accuracy of pest and disease detection. Modern computer vision and artificial intelligence technologies can monitor crop health in real-time through remote sensing devices and sensors, accurately identify pests and diseases, and reduce reliance on manual labor.
Precision agriculture uses technology to achieve efficient resource utilization, maximizing crop yields while minimizing environmental pollution. By dynamically adjusting resource allocation, optimizing planting plans, reducing pesticide and fertilizer use, and lowering carbon footprints, precision agriculture promotes sustainable agricultural practices. However, the implementation of precision agriculture depends on IoT sensors, remote sensing technology, and big data analysis to monitor soil, climate, and crop health factors in real-time, thereby optimizing agricultural decision-making.
The application of remote sensing technology in agriculture is widespread, particularly in crop health monitoring and large-scale farmland surveillance. As shown in Figure 1, mobile platforms collect weed images. While remote sensing images can provide real-time data, factors such as weather and climate changes affect image quality, necessitating improvements in clarity, resolution, and noise reduction in existing technologies. Additionally, efficiently processing and analyzing large-scale remote sensing data to extract useful information remains an urgent issue.
Figure 1: Weed image acquisition by a mobile platform with the OpenWeedGUI in two different fields. [33]
Therefore, there is an urgent need for the adoption of automated and intelligent technologies, such as deep learning-based crop pest and disease detection systems, precision agriculture decision support systems, and remote sensing image enhancement and analysis technologies. These technologies will help enhance production efficiency, accuracy, and sustainability, driving the transformation of agriculture from traditional models to intelligent and precision-oriented development.
@@ 2.2 Potential applications of diffusion models in agriculture
Diffusion models show great promise in agricultural applications, particularly in crop monitoring, pest and disease identification, and remote sensing image processing. The agricultural sector often faces challenges such as data scarcity and sample imbalance, but diffusion models can effectively address these issues due to their high-quality generation capabilities.
First, diffusion models have significant potential for pest and disease identification in crops. Traditional pest and disease detection methods rely on manual inspections, which are inefficient and prone to misdiagnosis, especially during the early stages of disease onset. Diffusion models can generate high-quality images of pests and diseases, particularly in the early stages when symptoms are not yet evident, providing clear diagnostic information to enhance the accuracy and efficiency of agricultural monitoring systems.
Second, in crop growth prediction, diffusion models can generate simulated crop growth data, as shown in Figure 2. By combining environmental factors such as soil moisture and climate change, they assist agricultural experts in predicting the health status and yield of crops at different growth stages, providing valuable support for scientific planting decisions.
Figure 2: Proposed pipeline to train Stable Diffusion to generate artificial images.[34]
Additionally, diffusion models show significant potential in enhancing and denoising remote sensing images. Remote sensing technology provides crucial data for agricultural monitoring, but image quality is often limited by weather and environmental conditions. Diffusion models can effectively improve the quality of remote sensing images, enabling agricultural experts to obtain clearer farmland data and enhance the accuracy of tasks such as crop health monitoring, land use assessment, and climate change impact analysis.
@@ 2.3 Advantages and challenges of diffusion models in agriculture
The advantages of diffusion models in agriculture primarily lie in their image generation quality, training stability, and flexible conditional generation capabilities. Compared to traditional generative adversarial networks (GANs), diffusion models exhibit higher stability during training, avoiding common issues such as pattern collapse that are associated with GANs, and can generate more diverse and high-quality agricultural images.
First, diffusion models generate high-quality images through gradual denoising, with details and clarity particularly prominent in complex agricultural environments. Whether in pest and disease monitoring or crop growth prediction, diffusion models can generate image data rich in detail, closely aligned with actual agricultural needs, significantly enhancing the accuracy and reliability of image generation.
Second, diffusion models offer a clear advantage in training stability compared to GANs. GANs often encounter instability during training and are prone to getting stuck in local optima. Diffusion models ensure training stability while improving generation quality through their gradual denoising process, making them suitable for handling large-scale data generation tasks in agriculture.
Additionally, diffusion models support flexible conditional generation, enabling the creation of images tailored to specific needs based on crop types, pest and disease categories, or environmental conditions. This capability allows diffusion models to provide customized solutions in agriculture, assisting farmers and agricultural experts with more precise crop monitoring, resource allocation, and precision agriculture decision-making.
Although diffusion models demonstrate significant advantages in agricultural applications, they still face challenges such as computational efficiency and data quality. Future research should focus on improving the computational efficiency of these models, enriching agricultural datasets, and exploring multimodal data fusion techniques to further enhance the application potential of diffusion models in agriculture.

### v3

@@ Abstract
With the global population increasing and arable land resources becoming increasingly limited, smart and precision agriculture have emerged as essential directions for sustainable agricultural development. Artificial intelligence (AI), particularly deep learning models, has been widely adopted in applications such as crop monitoring, pest detection, and yield prediction. Among recent generative models, diffusion models have demonstrated considerable potential in agricultural image processing, data augmentation, and remote sensing analysis. Compared to traditional generative adversarial networks (GANs), diffusion models exhibit greater training stability and superior image generation quality, effectively addressing challenges such as limited annotated datasets and imbalanced sample distributions in agricultural scenarios. This paper reviews recent advancements in the application of diffusion models within agriculture, focusing on their roles in crop disease and pest detection, remote sensing image enhancement, crop growth prediction, and agricultural resource management. Empirical studies show that diffusion models significantly enhance the performance of downstream models by improving accuracy, robustness, and generalization in tasks involving image synthesis, augmentation, and denoising under complex environmental conditions. Despite ongoing challenges in computational efficiency and domain generalization, diffusion models are expected to play an increasingly important role in the future of intelligent agriculture. As the technology continues to evolve, it holds substantial promise for addressing pressing global issues in food security and environmental sustainability.
@@
Diffusion Model, Smart Agriculture, Pest and Disease Detection, Crop Monitoring
@@ 1 Introduction
With the continuous growth of the global population and the increasing scarcity of arable land resources, traditional agriculture is facing unprecedented challenges. To improve production efficiency, ensure food security, and achieve sustainable development, countries around the world are actively promoting the modernization of agriculture. In this context, smart agriculture and precision farming—key pillars of agricultural modernization—are gradually becoming mainstream directions for future agricultural development.
@@
Smart agriculture integrates modern information technologies such as artificial intelligence (AI), the Internet of Things (IoT), big data, and remote sensing to enable dynamic sensing and data-driven decision-making for key factors including crop growth environments, equipment operating status, soil moisture, and climate conditions. Through precise fertilization, targeted irrigation, and intelligent pest and disease identification, smart agriculture significantly improves resource utilization while enhancing the intelligence, automation, and sustainability of agricultural production. As AI technologies continue to evolve, deep learning models—particularly image generation models—have become a focal point in agricultural intelligence research. Among them, diffusion models, a novel class of generative models, have shown remarkable performance in image generation [1, 2, 3, 4, 5, 6, 7], restoration [1, 3, 4, 8, 9, 10], segmentation [11, 12, 13, 14], and enhancement [15]. These models have achieved notable success in fields such as medicine [11, 13], industrial inspection [16], and artistic creation. Introducing diffusion models into agricultural scenarios offers promising solutions to challenges such as limited datasets, imbalanced samples, and insufficient data diversity.
This paper aims to systematically review the current state and typical applications of diffusion models in agriculture, thereby promoting the integration of AI and modern farming, and accelerating the practical implementation of agricultural intelligence. In this process, deep learning—known for its powerful feature extraction and pattern recognition capabilities—has become a cornerstone technology for agricultural informatization and automation. Unlike traditional machine learning, deep learning can automatically learn complex nonlinear relationships in an end-to-end fashion, significantly enhancing the analysis of images, time series, and multi-source data. Its applications span various subfields including crop identification, pest and disease detection [17, 18], yield prediction [19], and environmental monitoring [20, 21, 22].
@@
In image processing, convolutional neural networks (CNNs) [23] are widely used for tasks such as crop classification [24], weed identification, and leaf disease diagnosis [17, 18], by leveraging high-resolution images to detect fine-grained differences in agricultural fields. In remote sensing, deep neural networks effectively process data such as radar [25, 26] and multispectral imagery [27, 28, 29], enabling automated identification and analysis of crop types, planting areas, growth stages, and moisture content. For time-series prediction, recurrent neural networks (RNNs) and their variants such as long short-term memory (LSTM) networks are employed to model agricultural time-series data (e.g., soil moisture, meteorological parameters [22], and crop growth indicators), thus supporting precise fertilization and irrigation decisions.
@@
In addition, generative models like generative adversarial networks (GANs) [15, 30] have been increasingly adopted for agricultural data synthesis and augmentation, addressing data scarcity and sample imbalance. However, GANs suffer from training instability and mode collapse, which limit their broader adoption. In contrast, diffusion models offer greater training stability and higher-quality generation, making them increasingly valuable for agricultural applications.
@@
Diffusion models are a class of deep generative models that have recently made significant breakthroughs in generative modeling. Their foundation lies in the mathematical modeling of diffusion processes from non-equilibrium thermodynamics. Unlike GANs [15, 30] and variational autoencoders (VAEs) [31, 32], diffusion models generate high-quality samples by gradually adding noise to the data during training (forward process) and then reconstructing the data via learned denoising steps (reverse process). Since their introduction, diffusion models have undergone rapid development and achieved notable success across multiple domains. The denoising diffusion probabilistic model (DDPM), introduced by Ho et al. in 2020, marked the first large-scale application of diffusion models in deep learning, addressing key issues in training instability. DDIM further improved generation speed through a non-Markovian reverse process. Score-based generative models (SGMs) introduced new generation mechanisms by learning data score functions, mitigating some limitations of previous methods. Conditional diffusion models enable flexible conditional generation tasks such as text-to-image generation. Latent diffusion models (LDMs), proposed in 2022, further enhanced generation efficiency by operating in latent space, reducing computational cost and enabling high-resolution generation. The release of Stable Diffusion marked a milestone, offering powerful open-source tools for text-to-image generation.
Despite their strengths in generation quality and stability, diffusion models still face challenges including high computational demands, reliance on data quality, and generalization limitations. Future research is expected to improve efficiency, enable multimodal fusion, and adapt these models to low-resource environments, thereby expanding their practical applications and driving the evolution of generative AI in agriculture.
Figure 1: Structural overview of this review on diffusion models in agriculture.
The main contributions of this paper are summarized as follows:
  1. 1.
    Comprehensive and Structured Survey: We present the first comprehensive and structured review of diffusion models in smart agriculture, systematically categorizing their applications into seven key domains—namely, image classification, image generation, data augmentation, remote sensing and hyperspectral reconstruction, pest and disease diagnosis, multimodal fusion, and small-sample data generation. This categorization (illustrated in Figure 1) provides a clear and unified framework for understanding the diverse roles of diffusion models in agricultural intelligence.
  2. 2.
    Empirical Benchmarking and Practical Validation: Unlike prior theoretical surveys, we conduct small-scale empirical studies to benchmark diffusion models (such as DDPM and Stable Diffusion) against traditional GANs in typical agricultural tasks. Notably, we validate the effectiveness of diffusion models in addressing long-tail distribution problems through synthetic data augmentation. Our results demonstrate superior accuracy, robustness, and generalization in challenging agricultural scenarios.
  3. 3.
    In-depth Analysis of Model Variants and Technical Advancements: We detail the theoretical foundations, architectural evolutions, and application-specific adaptations of major diffusion model variants, including DDIM, Score-based Models, Conditional Diffusion Models, and Latent Diffusion Models. Their suitability and performance in agricultural applications are critically discussed, bridging algorithmic innovation with field-level deployment.
  4. 4.
    Identification of Key Challenges and Future Directions: We highlight current limitations—such as high computational cost, data dependency, and domain generalization—and propose future research directions involving lightweight model design, multimodal-guided generation, domain adaptation techniques, and real-time inference frameworks to facilitate broader adoption in practical agricultural environments.
  5. 5.
    Promotion of AI-Agriculture Integration: This work not only serves as a technical reference but also emphasizes the transformative potential of diffusion models in ensuring food security, promoting sustainable farming practices, and advancing environmental monitoring, thereby providing a roadmap for future interdisciplinary research at the intersection of artificial intelligence and smart agriculture.
The structure of this paper is organized as follows: Section 1 introduces the motivation, background, and significance of applying diffusion models in the context of modern agricultural development. Section 2 provides an overview of the agricultural sector and outlines the data characteristics, technical requirements, and practical challenges that justify the integration of diffusion models into agricultural applications. Section 3 reviews the theoretical foundations, development history, and representative variants of diffusion model technology, including key advances in generative modeling. Section 4 presents the specific applications of diffusion models in agricultural image analysis, such as classification, segmentation, data augmentation, image generation, and pest and disease detection. Section 5 reports experimental results and comparative analyses, highlighting the performance of diffusion models in various agricultural datasets. Section 6 discusses the current limitations and open challenges of diffusion models in agricultural contexts and proposes potential solutions and future research directions. Finally, Section 7 summarizes the key findings and contributions of the study, and outlines the prospects for diffusion model development in smart and precision agriculture.
@@ 2.1 Major Challenges and Technological Needs Facing Agriculture
Agriculture today faces several critical challenges, including crop health and pest management, precision resource allocation, and the efficient processing of remote sensing imagery. Pests and diseases are among the most significant threats to crop productivity. Traditional manual inspection methods are often inefficient and prone to misdiagnosis, especially in the early stages of disease development. As the scale of agricultural operations continues to grow, intelligent technologies have become essential for enhancing the efficiency and accuracy of pest and disease detection. With the integration of modern computer vision and artificial intelligence technologies, crop health can now be monitored in real time using remote sensing devices and sensor networks, enabling accurate identification of pests and diseases while reducing reliance on manual labor.
Precision agriculture leverages advanced technologies to optimize resource utilization, thereby maximizing yields while minimizing environmental impacts. It enables dynamic adjustments to planting strategies, fertilizer and pesticide application, and irrigation practices, contributing to reduced carbon emissions and sustainable agricultural development. However, implementing precision agriculture relies heavily on IoT sensors, remote sensing technologies, and big data analytics to monitor soil, weather, and crop health conditions in real time, thereby supporting data-driven agricultural decision-making.
Remote sensing technology plays an increasingly prominent role in agricultural applications, especially for large-scale monitoring of crop health and farmland status. As illustrated in Figure 2, mobile platforms can be used to collect weed imagery. Despite the advantages of real-time data acquisition, remote sensing images are often affected by weather and lighting conditions, leading to degraded quality. Enhancing resolution, clarity, and noise reduction in such images remains a pressing technological challenge. Moreover, efficiently processing and analyzing large-scale remote sensing datasets to extract meaningful agricultural insights is another urgent issue.
Figure 2: Weed image acquisition by a mobile platform with OpenWeedGUI in two different fields [33].
To address these challenges, there is a growing need to integrate intelligent and automated technologies, such as deep learning-based pest and disease detection systems, decision support systems for precision agriculture, and advanced image enhancement techniques for remote sensing. These innovations offer the potential to significantly improve agricultural productivity, decision-making accuracy, and sustainability, facilitating the transition from traditional farming to intelligent, precision-driven agriculture.
@@ 2.2 Potential Applications of Diffusion Models in Agriculture
Diffusion models have shown considerable promise for agricultural applications, particularly in crop monitoring, pest and disease detection, and remote sensing image analysis. In scenarios characterized by data scarcity and sample imbalance, diffusion models are especially valuable due to their ability to generate high-quality synthetic data.
One of the most promising applications lies in the early detection of crop pests and diseases. Traditional detection approaches rely heavily on manual inspection, which is inefficient and often fails to identify subtle or early-stage symptoms. Diffusion models can generate high-resolution images of pests and diseases, providing supplementary data that enhances the accuracy and efficiency of automated detection systems.
In crop growth prediction, diffusion models can generate synthetic growth data under various environmental conditions, as illustrated in Figure 3. By incorporating variables such as soil moisture and climatic changes, these models help forecast crop health and yield across different growth stages, thus supporting scientific planting and management decisions.
Figure 3: Proposed pipeline for training Stable Diffusion to generate artificial images [34].
Diffusion models also offer considerable potential in remote sensing image enhancement and denoising. Although remote sensing provides valuable data for agricultural monitoring, image quality is frequently compromised by environmental noise. Diffusion-based enhancement techniques can significantly improve the clarity and resolution of these images, facilitating more accurate land-use assessment, crop condition analysis, and climate impact monitoring.
@@ 2.3 Advantages and Challenges of Diffusion Models in Agriculture
The primary advantages of diffusion models in agricultural contexts include their high-quality image generation, superior training stability, and support for flexible conditional generation. Compared to traditional generative adversarial networks (GANs), diffusion models are more stable during training and less susceptible to issues such as mode collapse, making them well-suited for generating diverse and realistic agricultural imagery.
First, diffusion models achieve high image quality by progressively refining noisy data, which is particularly beneficial in complex agricultural scenarios. Whether for pest detection or crop growth simulation, these models can generate detailed, high-fidelity images that reflect real-world agricultural conditions, thereby improving the reliability and interpretability of data-driven decision-making.
Second, diffusion models offer significantly better training stability than GANs. While GANs often suffer from convergence instability and mode dropping, diffusion models employ a gradual denoising process that enhances robustness and reproducibility. This characteristic makes them ideal for large-scale image generation tasks in agriculture, where stable training is critical.
Third, diffusion models support conditional generation, enabling them to produce images tailored to specific requirements such as crop type, pest species, or environmental condition. This flexibility enables the development of customizable agricultural intelligence solutions that support more precise monitoring, decision-making, and resource optimization.
Despite these advantages, diffusion models still face several challenges. High computational demands, long inference times, and dependence on high-quality training data can limit their scalability in real-world agricultural settings. Future research should focus on improving computational efficiency, enriching agricultural datasets, and exploring multimodal data fusion techniques to further expand the applicability of diffusion models in agriculture.